# How to avoid machine learning pitfalls: a guide for academic researchers

### v4

@@ 4.4 Do evaluate a model multiple times
Many ML models are unstable. That is, if you train them multiple times, or if you make small changes to the training data, then their performance varies significantly. This means that a single evaluation of a model can be unreliable, and may either underestimate or overestimate the model’s true potential. For this reason, it is common to carry out multiple evaluations. There are numerous ways of doing this, but most involve training the model multiple times using different subsets of the training data. Cross-validation (CV) is particularly popular, and comes in numerous flavours (Arlot et al., 2010), most of which involve splitting the data into a number of folds. When doing CV, it is important to be aware of any dependencies within the data and take these into account. Failure to do so can result in data leakage. For instance, in medical datasets, it is commonplace to have multiple data points for a single subject; to avoid data leakage, these should be kept together within the same fold. Time series data is particularly problematic for CV; see Don’t ignore temporal dependencies in time series data for a discussion of how to handle this. If you’re carrying out hyperparameter optimisation, then you should use nested cross-validation (also known as double cross-validation), which uses an extra loop inside the main cross-validation loop to avoid overfitting the test folds. If some of your data classes are small, then you may need to do stratification, which ensures each class is adequately represented in each fold.
@@ 4.5 Do save some data to evaluate your final model instance
I’ve used the term model quite loosely, but there is an important distinction between evaluating the potential of a general model (e.g. how well a neural network can solve your problem), and the performance of a particular model instance (e.g. a specific neural network produced by one run of back-propagation). Cross-validation is good at the former, but it’s less useful for the latter. Say, for instance, that you carried out ten-fold cross-validation. This would result in ten model instances. Say you then select the instance with the highest test fold score as the model which you will use in practice [^4] Another approach sometimes used in this situation is to ensemble the models; see Do combine models (carefully). However, this could have downsides in terms of model complexity and inference time.. How do you report its performance? Well, you might think that its test fold score is a reliable measure of its performance, but it probably isn’t. First, the amount of data in a single fold is relatively small. Second, the instance with the highest score could well be the one with the easiest test fold, so the evaluation data it contains may not be representative. Consequently, the only way of getting a reliable estimate of the model instance’s generality may be to use another test set. This is also true in situations where the independence of the existing test set may have been compromised, e.g. by using it more than once (see Do avoid sequential overfitting). So, if you have enough data, it’s better to keep some aside and only use it once to provide an unbiased estimate of the final selected model instance.
@@ 4.6 Do choose metrics carefully
Figure 5: See Do choose metrics carefully.
Be careful which metrics you use to evaluate your ML models. For instance, in the case of classification models, the most commonly used metric is accuracy, which is the proportion of samples in the data set that were correctly classified by the model. This works fine if your classes are balanced, i.e. if each class is represented by a similar number of samples within the data set. But many data sets are not balanced, and in this case accuracy can be a very misleading metric. Consider, for example, a data set in which 90% of the samples represent one class, and 10% of the samples represent another class. A binary classifier which always outputs the first class, regardless of its input, would have an accuracy of 90%, despite being completely useless (see Figure 4.6). In this kind of situation, it would be preferable to use a metric such as $F_{1}$ score, Cohen’s kappa coefficient ($\kappa$) or Matthews Correlation Coefficient (MCC), all of which are relatively insensitive to class size imbalance. For a broader review of methods for dealing with imbalanced data, see Haixiang et al. (2017). There are also various pitfalls associated with regression metrics, particularly within the context of time series forecasting; see Hewamalage et al. (2023) for a discussion of these. A well-known example is relying only on the RMSE, which (a bit like accuracy) is susceptible to assigning high value to models which always predict no change. Also see Do report performance in multiple ways.
@@ 4.7 Don’t ignore temporal dependencies in time series data
Time series data is unlike many other kinds of data in that the order of the data points is important. Many of the pitfalls in handling time series data are a result of ignoring this fact. Most notably, time series data are subject to a particular kind of data leakage (see Don’t allow test data to leak into the training process) known as look ahead bias. This occurs when some or all of the data points used to train the model occur later in the time series than those used to test the model. In effect, this can allow knowledge of the future to leak into training, and this can then bias the test performance. A situation where this commonly occurs is when standard cross-validation (see Do evaluate a model multiple times) is applied to time series data, since it results in the training folds in all but one of the cross-validation iterations containing data that is in the future relative to the test fold. This can be avoided by using special forms of cross-validation that respect temporal dependencies, such as blocked cross-validation, though whether this is necessary depends to some extent on the nature of the time series data, e.g. whether it is stationary or non-stationary. See Cerqueira et al. (2020) and Wang and Ruf (2022) for more on this. Look ahead bias can also result from carrying out data-dependent preprocessing operations before splitting off the test data; see Fig. 6 for a simple example of this, but also see Do be careful where and how you do feature selection.
@@ 5.3 Do correct for multiple comparisons
Things get a bit more complicated when you want to use statistical tests to compare more than two models, since doing multiple pairwise tests is a bit like using the test set multiple times — it can lead to overly-optimistic interpretations of significance. Basically, each time you carry out a comparison between two models using a statistical test, there’s a probability that it will discover significant differences where there aren’t any. This is represented by the confidence level of the test, usually set at 95%: meaning that 1 in 20 times it will give you a false positive. For a single comparison, this may be a level of uncertainty you can live with. However, it accumulates. That is, if you do 20 pairwise tests with a confidence level of 95%, one of them is likely to give you the wrong answer. This is known as the multiplicity effect, and is an example of a broader issue in data science known (at least when done intentionally) as data dredging or p-hacking — see Head et al. (2015). To address this problem, you can apply a correction for multiple tests. The most common approach is the Bonferroni correction, a very simple method that lowers the significance threshold based on the number of tests that are being carried out; see Salzberg (1997) for a gentle introduction. However, there are numerous other approaches, and there is also some debate about when and where these corrections should be applied; for an accessible overview, see Streiner (2015).
@@ 5.4 Don’t always believe results from community benchmarks
In certain problem domains, it has become commonplace to use benchmark data sets to evaluate new ML models. The idea is that, because everyone is using the same data to train and test their models, then comparisons will be more transparent. Unfortunately this approach has some major drawbacks. First, if access to the test set is unrestricted, then you can’t assume that people haven’t used it as part of the training process. This is known as “developing to the test set”, and leads to results that are heavily over-optimistic. A more subtle problem is that, even if everyone who uses the data only uses the test set once, collectively the test set is being used many times by the community. In effect, by comparing lots of models on the same test set, it becomes increasingly likely that the best model just happens to over-fit the test set, and doesn’t necessarily generalise any better than the other models (see Do correct for multiple comparisons and Do avoid sequential overfitting). For these, and other reasons, you should be careful how much you read into results from a benchmark data set, and don’t assume that a small increase in performance is significant. This is particularly the case where foundation models (see Do keep up with progress in deep learning (and its pitfalls)) are used, since it is possible that their training data included the test sets from community benchmarks. See Paullada et al. (2020) for a wider discussion of issues surrounding the use of shared datasets. Also see Do report performance in multiple ways.
@@ 6.5 Do look at your models
Trained models contain a lot of useful information. Unfortunately many authors just report the performance metrics of a trained model, without giving any insight into what it actually learnt. Remember that the aim of research is not to get a slightly higher accuracy than everyone else. Rather, it’s to generate knowledge and understanding and share this with the research community. If you can do this, then you’re much more likely to get a decent publication out of your work. So, do look inside your models and do try to understand how they reach a decision. For relatively simple models like decision trees, it can also be beneficial to provide visualisations of your models, and most libraries have functions that will do this for you. For complex models, like deep neural networks, consider using explainable AI (XAI) techniques to extract knowledge (surveyed in Li et al. (2020) and Angelov et al. (2021)); they’re unlikely to tell you exactly what the model is doing (see Don’t assume deep learning will be the best approach), but they may give you some useful insights.
@@ 6.6 Do use a machine learning checklist
This guide aims to give an appreciation of the main things that can go wrong during machine learning, plus some guidance on how to avoid these things going wrong. Checklists, on the other hand, are designed to take you more formally through the ML pipeline and encourage you to document (and more importantly, think about) how your implementation decisions support a meaningful outcome. In some domains, e.g. certain fields of medicine, it is compulsory to complete a checklist before submitting a paper for publication. However, beyond their quality assurance role, checklists are arguably most useful at the start of a study when it comes to planning an ML pipeline. Since I’m one of the authors, I’d particularly encourage you to look at REFORMS (Kapoor et al., 2023), which is a combined checklist and set of consensus-based recommendations for doing ML-based science (although much of it is also applicable to ML practice more generally). However, other, more domain-specific, checklists are available.
@@ 7 Final thoughts
This document doesn’t tell you everything you need to know, the lessons sometimes have no firm conclusions, and some of the things I’ve told you might be wrong, or at least debatable. This, I’m afraid, is the nature of research. The theory of how to do ML almost always lags behind the practice, academics will always disagree about the best ways of doing things, and what we think is correct today may not be correct tomorrow. Therefore, you have to approach ML in much the same way you would any other aspect of research: with an open mind, a willingness to keep up with recent developments, and the humility to accept you don’t know everything.
8 Acknowledgements
Many thanks to everyone who gave me feedback on the draft manuscript, and to everyone who has since sent me suggestions for revisions and new content.
9 Changes
Changes from v3
Added Do use a machine learning checklist and Do think about how and where you will use data. Rewrote Do evaluate a model multiple times. Revised/extended (and renamed) Do keep up with progress in deep learning (and its pitfalls), Do be careful where and how you do feature selection, Do avoid sequential overfitting, Do choose metrics carefully and Do combine models (carefully). Extended Do use an appropriate test set. Added references (Aguinis et al., 2021; Hewamalage et al., 2023; Hosseini et al., 2020; Kapoor et al., 2023; Li et al., 2023; Molnar et al., 2020; Roberts et al., 2021; Zhou et al., 2023). Various tweaks elsewhere.
Changes from v2
Added illustrations. Added Do avoid learning spurious correlations, Don’t ignore temporal dependencies in time series data and Do keep up with progress in deep learning (and its pitfalls). Added references (Cerqueira et al., 2020; Gibney, 2022; Grinsztajn et al., 2022; Han et al., 2021; Iwana and Uchida, 2021; Kapoor and Narayanan, 2022; Khan et al., 2022; Li et al., 2021; Liao et al., 2021; Lin et al., 2022; Liu et al., 2021; Schmidhuber, 2015; Shankar et al., 2022; Zhang et al., 2021). Various tweaks elsewhere.
Changes from v1
Added Don’t do data augmentation before splitting your data and Don’t assume deep learning will be the best approach. Rewrote Don’t use inappropriate models. Expanded Don’t allow test data to leak into the training process, Do be careful when reporting statistical significance and Do be transparent. Added references (Angelov et al., 2021; Benavoli et al., 2017; Bischl et al., 2021; Chen et al., 2020; Vandewiele et al., 2021; Wang and Ruf, 2022). Various tweaks elsewhere.

### v5

@@ 4.4 Do evaluate a model multiple times
Many ML models are stochastic or unstable. That is, if you train them multiple times, or if you make small changes to the training data, then their performance varies significantly. The same is true of using LLMs at inference time. This means that a single evaluation of a model can be unreliable, and may either underestimate or overestimate the model’s true potential. For this reason, it is common to carry out multiple evaluations. At training time, there are numerous ways of doing this. For stochastic models, the simplest is to train the same model multiple times using different random seeds and then look at the average performance. A more robust approach is to also vary the data for each model trained. Cross-validation (CV) is a particularly popular way of doing this, and comes in numerous flavours (Arlot et al., 2010), most of which involve splitting the data into a number of folds. When doing CV, it is important to be aware of any dependencies within the data and take these into account. Failure to do so can result in data leakage. For instance, in medical datasets, it is commonplace to have multiple data points for a single subject; to avoid data leakage, these should be kept together within the same fold. Time series data is particularly problematic for CV; see Don’t ignore temporal dependencies in time series data for a discussion of how to handle this. If you’re carrying out hyperparameter optimisation, then you should use nested cross-validation (also known as double cross-validation), which uses an extra loop inside the main cross-validation loop to avoid overfitting the test folds. If some of your data classes are small, then you may need to do stratification, which ensures each class is adequately represented in each fold. In addition to looking at average performance across multiple evaluations, it is also standard practice to provide some measure of spread or confidence, such as the standard deviation or the 95% confidence interval.
@@ 4.5 Do save some data to evaluate your final model instance
I’ve used the term model quite loosely, but there is an important distinction between evaluating the potential of a general model (e.g. how well a neural network can solve your problem), and the performance of a particular model instance (e.g. a specific neural network produced by one run of back-propagation). Cross-validation is good at the former, but it’s less useful for the latter. Say, for instance, that you carried out ten-fold cross-validation. This would result in ten model instances. Say you then select the instance with the highest test fold score as the model which you will use in practice. How do you report its performance? Well, you might think that its test fold score is a reliable measure of its performance, but it probably isn’t. First, the amount of data in a single fold is relatively small. Second, the instance with the highest score could well be the one with the easiest test fold, so the evaluation data it contains may not be representative. Consequently, the only way of getting a reliable estimate of the model instance’s generality may be to use another test set. This is also true in situations where the independence of the existing test set may have been compromised, e.g. by using it more than once (see Do avoid sequential overfitting). So, if you have enough data, it’s better to keep some aside and only use it once to provide an unbiased estimate of the final selected model instance. However, it’s worth noting one other option when using cross-validation, and that is to ensemble the model instances (see Do combine models (carefully)). The resulting ensemble will have performance in line with the average as measured through cross-validation, so another test set is not required to measure its performance. On the downside, it will likely have poorer inference time, efficiency and interpretability than a single model instance, so this approach is generally only worth considering if you have very little data.
@@ 4.6 Do choose metrics carefully
Figure 6: See Do choose metrics carefully. The problem with using accuracy as a performance metric on imbalanced data. Here, a dummy model which always predicts the same class label has an accuracy of 50% or 90% depending on the distribution of class labels within the data.
Be careful which metrics you use to evaluate your ML models. For instance, in the case of classification models, the most commonly used metric is accuracy, which is the proportion of samples in the data set that were correctly classified by the model. This works fine if your classes are balanced, i.e. if each class is represented by a similar number of samples within the data set. But many data sets are not balanced, and in this case accuracy can be a very misleading metric. Consider, for example, a data set in which 90% of the samples represent one class, and 10% of the samples represent another class. A binary classifier which always outputs the first class, regardless of its input, would have an accuracy of 90%, despite being completely useless (see Figure 6). In this kind of situation, it would be preferable to use a metric such as $F_{1}$ score, Cohen’s kappa coefficient ($\kappa$) or Matthews Correlation Coefficient (MCC), all of which are relatively insensitive to class size imbalance. For a broader review of methods for dealing with imbalanced data, see Haixiang et al. (2017). There are also various pitfalls associated with regression metrics, particularly within the context of time series forecasting; see Hewamalage et al. (2023) for a discussion of these. A well-known example is relying only on the RMSE, which (a bit like accuracy) is susceptible to assigning high value to models which always predict no change. Also see Do report performance in multiple ways.
4.7 Do consider model fairness
Overall performance metrics are not the only important measures of how good a model is. If a model is to be deployed within the real world, another important measure is fairness. There are various definitions of fairness, but in a nutshell, it’s about making sure that the model doesn’t treat its human subjects unequally with regard to characteristics such as gender, ethnicity, income or personal politics. This is also referred to as algorithmic bias, and there are many examples of models being biased towards or against particular groups of people. A common source of unfairness is using an unrepresentative dataset to train an ML model. For instance, if a medical diagnosis model is trained on data from a single country, then the data may be biased towards the majority ethnicity, and the model may not operate fairly when exposed to users from other ethnicities. However, unfairness can also come from other sources, including subconscious bias during data preparation and the inductive biases of the model. Regardless of the source, it is important to understand any resulting biases, and ideally take steps to mitigate against them (e.g. applying data augmentation to minority samples — see Do make sure you have enough data). There are many different fairness metrics, so part of the puzzle is working out which are most relevant to your modelling context; see Caton and Haas (2024) for a review.
@@ 4.8 Don’t ignore temporal dependencies in time series data
Time series data is unlike many other kinds of data in that the order of the data points is important. Many of the pitfalls in handling time series data are a result of ignoring this fact. Most notably, time series data are subject to a particular kind of data leakage (see Don’t allow test data to leak into the training process) known as look ahead bias. This occurs when some or all of the data points used to train the model occur later in the time series than those used to test the model. In effect, this can allow knowledge of the future to leak into training, and this can then bias the test performance. A situation where this commonly occurs is when standard cross-validation (see Do evaluate a model multiple times) is applied to time series data, since it results in the training folds in all but one of the cross-validation iterations containing data that is in the future relative to the test fold. This can be avoided by using special forms of cross-validation that respect temporal dependencies, such as blocked cross-validation, though whether this is necessary depends to some extent on the nature of the time series data, e.g. whether it is stationary or non-stationary. See Cerqueira et al. (2020) and Wang and Ruf (2022) for more on this. Look ahead bias can also result from carrying out data-dependent preprocessing operations before splitting off the test data; see Figure 7 for a simple example of this, but also see Do be careful where and how you do feature selection.
@@ 5.1 Don’t assume a bigger number means a better model
5.2 Do use meaningful baselines
When introducing a new modelling approach, it is essential to compare against established approaches. These are commonly referred to as baseline models, or just baselines. It is important that these baselines are selected so that they provide a meaningful basis for comparison. Baselines are often simpler than the new approach, and are chosen to demonstrate that any complexity in the new model is necessary. For example, if you’re extending model X, then it makes sense to use model X as a baseline. However, it also makes sense to use other simpler models. For instance, if you’re developing a deep learning approach that uses tabular data, then you should also compare against simpler models like decision trees and SVMs to show that a more complex approach is justified. If you’re solving a regression problem, then you should also consider using simple baselines like logistic regression. The simplest baselines are known as naïve baselines and are used to show that your model is not doing something trivial. An illustrative example of why these are necessary is described in Hewamalage et al. (2023), where a complex transformer model designed for time series forecasting is shown to perform worse than a naïve baseline that always forecasts the next value in a time series to be the same as the previous value. This kind of naïve baseline, in which there is no real decision making process, is also known as a dummy model. Another example is a classifier that always outputs the most frequent class label (as described in Do choose metrics carefully). In addition to simple baselines, it’s also important to compare against state-of-the-art (SOTA) models. Otherwise you may be asked something like “Why are you extending model X when model Y is known to be better than model X?”
@@ 5.4 Do correct for multiple comparisons
Things get a bit more complicated when you want to use statistical tests to compare more than two models, since doing multiple pairwise tests is a bit like using the test set multiple times — it can lead to overly-optimistic interpretations of significance. Basically, each time you carry out a comparison between two models using a statistical test, there’s a probability that it will discover significant differences where there aren’t any. This is represented by the confidence level of the test, usually set at 95%: meaning that 1 in 20 times it will give you a false positive. For a single comparison, this may be a level of uncertainty you can live with. However, it accumulates. That is, if you do 20 pairwise tests with a confidence level of 95%, one of them is likely to give you the wrong answer. This is known as the multiplicity effect, and is an example of a broader issue in data science known (at least when done intentionally) as data dredging or p-hacking — see Stefan and Schönbrodt (2023). To address this problem, you can apply a correction for multiple tests. The most common approach is the Bonferroni correction, a very simple method that lowers the significance threshold based on the number of tests that are being carried out; see Salzberg (1997) for a gentle introduction. However, there are numerous other approaches, and there is also some debate about when and where these corrections should be applied; for an accessible overview, see Streiner (2015).
@@ 5.5 Don’t always believe results from community benchmarks
In certain problem domains, it has become commonplace to use benchmark data sets to evaluate new ML models. The idea is that, because everyone is using the same data to train and test their models, then comparisons will be more transparent. Unfortunately this approach has some major drawbacks. First, if access to the test set is unrestricted, then you can’t assume that people haven’t used it as part of the training process. This is known as “training to the test set”, and leads to results that are heavily over-optimistic. A more subtle problem is that, even if everyone who uses the data only uses the test set once, collectively the test set is being used many times by the community. In effect, by comparing lots of models on the same test set, it becomes increasingly likely that the best model just happens to over-fit the test set, and doesn’t necessarily generalise any better than the other models (see Do correct for multiple comparisons and Do avoid sequential overfitting). For these, and other reasons, you should be careful how much you read into results from a benchmark data set, and don’t assume that a small increase in performance is significant. This is particularly the case where foundation models (see Do keep up with progress in deep learning (and its pitfalls)) are used, since it is possible that their training data included the test sets from community benchmarks. See Paullada et al. (2021) for a wider discussion of issues surrounding the use of shared datasets. Also see Do report performance in multiple ways.
@@ 6.5 Do look at your models
Figure 8: See Do look at your models. Using saliency maps to analyse vision-based deep learning models. Imagine these two maps (in red) were generated for the image shown in the centre, for two different deep learning models trained on the kind of tank recognition data mentioned in Do avoid learning spurious correlations. Darker colours indicate features that are of greater importance to the model, so the model on the left (which predominantly focuses on the components of the tank) is likely to generalise much better than the one on the right (which predominantly focuses on the background of the image).
Trained models contain a lot of useful information. Unfortunately many authors just report the performance metrics of a trained model, without giving any insight into what it actually learnt. Remember that the aim of research is not to get a slightly higher accuracy than everyone else. Rather, it’s to generate knowledge and understanding and share this with the research community. If you can do this, then you’re much more likely to get a decent publication out of your work. So, do look inside your models and do try to understand how they reach a decision. For relatively simple models like decision trees, it can also be beneficial to provide visualisations of your models, and most libraries have functions that will do this for you. For more complex models, there are a range of explainable AI (XAI) techniques that can be used. Some of these are model-specific, and others are model-agnostic. Well-established examples of the latter are LIME and SHAP; both give insights into which features are important for a model. For CNNs and vision transformers, a common approach is to use saliency maps, which show the importance of different parts of an input image — see Figure 8 for an illustrative example. Grad-CAM is a popular technique for generating these, but there are plenty of other methods too. For non-vision transformers, a common approach is to visualise attention weights. See Dwivedi et al. (2023) for a survey of XAI techniques, and Ali et al. (2023) for a discussion of the limitations of current approaches. Whilst XAI techniques can give you useful insights into a model’s behaviour, it’s important to bear in mind that they are unlikely to tell you exactly what a model is doing. This is particularly the case for deep learning models (see Don’t assume deep learning will be the best approach), whose complexity makes their behaviour inherently difficult to analyse. For complex models, ablation studies Meyes et al. (2019) can also be useful. This involves successively removing parts of the model to see what is important, and can result in a simpler model which is more amenable to analysis.
@@ 6.6 Do use a machine learning checklist
This guide aims to give an appreciation of the main things that can go wrong during machine learning, plus some guidance on how to avoid these things going wrong. Checklists, on the other hand, are designed to take you more formally through the ML pipeline and encourage you to document (and more importantly, think about) how your implementation decisions support a meaningful outcome. In some domains, e.g. certain fields of medicine, it is compulsory to complete a checklist before submitting a paper for publication. However, beyond their quality assurance role, checklists are arguably most useful at the start of a study when it comes to planning an ML pipeline. Since I’m one of the authors, I’d particularly encourage you to look at REFORMS (Kapoor et al., 2024), which is a combined checklist and set of consensus-based recommendations for doing ML-based science (although much of it is also applicable to ML practice more generally). Other, more domain-specific, checklists are also available.
@@ 7 Final thoughts
ML is becoming an important part of people’s lives, yet the practice of ML is arguably in its infancy. There are many easy-to-make mistakes that can cause an ML model to appear to perform well, when in reality it does not. In turn, this has the potential to misinform when these models are published, and the potential to cause harm if these models are ever deployed. This guide describes the most common of these mistakes, and also touches upon more general issues of good practice in ML, such as fairness, transparency and the avoidance of bias. It also offers advice on avoiding these pitfalls. However, new threats continue to emerge as new approaches to ML are developed, and it is therefore important for users of ML to remain vigilant. This is the nature of a fast-moving research area — the theory of how to do ML almost always lags behind the practice, practitioners will always disagree about the best ways of doing things, and what we think is correct today may not be correct tomorrow.
You can find more on ML pitfalls at my Substack, Fetch Decode Execute.
Acknowledgements
Many thanks to everyone who gave me feedback on the draft manuscript, to everyone who has since sent me suggestions for revisions and new content, and to the editor and peer reviewers of the version published in Patterns.